Grade 12th passDiscuss with colleagues and IITians

A motor van weighing 4400 kg rounds a level curve of radius 200 m on an unbanked road at 60 km/hr. What should be minimum value of coefficient of friction to prevent skidding? At what angle the road should be banked for this velocity?

xavier
9 Years agoGrade 12th pass

1 Answer

Devansh Jatin Ponda
9 Years ago
Hi
I got the first part . But second i got a negative value that tan theta = -0.001. 
let c be coefficient of friction.
v= root(cgr)
so 50/3 = root( c *2000)
so 2500/9 = c*2000
so c= 5/ 36 = 0.138

Hope it is useful!!